OpenStudy (anonymous):

definite integral on interval [4,9] of 2+x/2sqrt(x) dx?

OpenStudy (anonymous):

\[\int\limits_{4}^{9}\frac{ 2+x }{ 2\sqrt{x}}dx\]

OpenStudy (raffle_snaffle):

well did you try. what was your solution?

OpenStudy (anonymous):

i dont know what to u-sub

OpenStudy (anonymous):

1.break in two integrals or 2.substitute \[\sqrt{x}=t,x=t^2,dx=2 t dt,\] when x=4,t=? similarly find t for x= 9 complete the problem.